OpenStudy (anonymous):

Determine the point of intersection between the line R = (4,-3,15)+ t(2,-3,5) and xy plane.

OpenStudy (anonymous):

The answer is (-2,6,0) But how?

OpenStudy (amistre64):

xy plane is when z=0 right?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

x = x0-2t y = y0+3t z = z0-5t x = 4-2t y = -3+3t z = 15-5t = = 0 when t=3 right?

OpenStudy (anonymous):

OH YES GOT IT ! OMG THANKYOU SO STUPID OF ME !

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

MEDAL FOR YOU :)